# RetailKLIP : Finetuning OpenCLIP backbone using metric learning on a single GPU for Zero-shot retail product image classification

### 16 Dec 2023 v1

@@ 1 INTRODUCTION
Our work has two components : 1. Creating a large dataset which can be used for finetuning a CLIP backbone for retail product images. 2. Proposing a learning rate strategy and loss which solve for erratic finetuning of CLIP backbones and address imbalanced large retail datasets.
@@ 2 RELATED WORK
Deep Metric Learning [Musgrave et al., 2020] techniques can be used to learn encoders which generate discriminative embeddings. Given that retail datasets are very imabalanced, we choose metric learning techniques to finetune the CLIP model to the retail domain. ArcFace loss, used generally in long tailed facial recognition tasks is used to finetune CLIP [Deng et al., 2019]. To our knowledge, this is the first time metric learning has been used to finetune a LVM to a new domain.
@@ 3 DATASETS
We would like to list the datasets used for the experiments in our work. To finetune CLIP image encoder to retail product image recognition, we use an inhouse (not publicly available) dataset called RP6K with 6500 retail products. This has over 1 Million images of retail products with their class tags. Just like real world, the cardinality of these products in datasets in in power law with some products having upto 4000 images, while many having lesser than 10 images. We finetune CLIP image encoder on this dataset using a novel mixture of techniques which yields a model that can be used for Zero shot classification on other datasets.
Grozi-120 is the first dataset used for evaluation [Merler et al., 2007]. It is a one shot classification dataset with one train image per product, which is packshot like and real world images in the test set. It has 120 products.
@@
RP2K [Peng et al., 2020] is a dataset with 2388 retail products, with substancial number of both train and test images for each product. However, we still use the dataset in a few shot setting like Grozi120 and CAPG-GP. That is, while the entire test set will be used to test the performance of Zero Shot classification, only one image per class will be used to calculate class representations for classification.
@@ 4 METHOD
Our work involves finetuning OpenCLIP image encoder on a retail product image dataset RP6K such that resultant image encoder can yield embeddings of product images useful for zero shot retail product image classification. The image encoder after finetuning with RP6K is referred to as RetailKLIP. For zero shot classification, one image per product is taken and augmented and then passed to RetailKLIP to get embeddings for that product. This process is repeated for all products to create a set of products’ embeddings. To test an unseen image, its image embedding created using RetailKLIP is compared to product embeddings created earlier and the product with closest embedding is considered to be the output of RetailKLIP.
@@ 4.1 Finetuning OpenCLIP to RetailKLIP
We use the ViT-L OpenCLIP pretrained model to obtain RetailKLIP through finetuning. Specific Architecture used is ViT-L/14 trained on LAION-2B available on OpenCLIP’s Github [Ilharco et al., 2021]. We take the trick of using AdamW optimizer to finetune CLIP from [Kumar et al., 2022]. We also use differing learning rate with depth like in [Dong et al., 2022], but instead of layerwise rate decay, we use block wise rate decay. That is, within each block of ViT there is a common learning rate, but learning rate varies across blocks. The last block starts with learning rate of 2e-4 and each previous block having a decreasing learning rate by a factor of 0.7. RP6K is a very imbalanced dataset, hence we use metric learning loss to finetune OpenCLIP. Specifically, ArcFace [Deng et al., 2019] implementation of PyTorch Metric Learning library [Musgrave et al., 2020] is used for the task. ArcFace is considered good for datasets with a long tail distribution.
@@ 4.2 Zero Shot classification using RetailKLIP
To use RetailKLIP as a Zero Shot classifier on a new dataset, one doesn’t need to finetune RetailKLIP. One image per product is taken from trainset of product and its augmentations are created, These augmentations of a product image are passed through RetailKLIP to get a set of embeddings for the product. Once we have embeddings for all products, we are ready to classify. An image of the test dataset is classified by using RetailKLIP to get its embedding and compare it with existing products’ embeddings. The nearest embedding from train set is taken to be the result of classification.
@@ 5.1 RP2K results
RP2K dataset paper proposes full finetuning ResNet and other backbones on their dataset one category at a time to get upto 95% accuracy in some categories of their dataset. For some other harder categories, the accuracy can be lower than 90%. However, due to lack of Kanji knowledge, our team has been unable to separate out categories and thus we have to use models on all categories combined, so this makes the problem harder for RetailKLIP. Also RetailKLIP takes just one image from train set to classify test images unlike full finetuning which uses the ample amount of images available. The accuracy for Zero Shot classification on RP2K is 87.7%. This is close to the accuracy full finetuning approach of [Peng et al., 2020] gets on the hardest categories.

### 14 Jan 2024 v2

@@ 1 INTRODUCTION
Our work has two components : 1. Creating a large dataset which can be used for finetuning a CLIP backbone for retail product images. 2. Proposing a learning rate strategy, class balancing approach and other finetuning components which solve for erratic finetuning of CLIP backbones and address imbalanced large retail datasets.
@@ 2 RELATED WORK
Deep Metric Learning [Musgrave et al., 2020] techniques can be used to learn encoders which generate discriminative embeddings. Unlike softmax based classification, metric learning approaches generalize better to openset recognition [Deng et al., 2019]. ArcFace loss, a metric learning loss function, used generally in long tailed facial recognition tasks is used to finetune CLIP [Deng et al., 2019]. To our knowledge, this is the first time metric learning has been used to finetune a LVM to a new domain.
Given that retail datasets are very imabalanced is another reason we use ArcFace loss to finetune the CLIP model to the retail domain. Previous works have shown that plain ArcFace handles imbalanced datasets better than softmax loss even when aided with data balancing techniques [Zhang and Deng, 2020] . We use ArcFace loss with a custom data balancing technique for our final results.
@@ 3 DATASETS
We would like to list the datasets used for the experiments in our work. To finetune CLIP image encoder to retail product image recognition, we use an inhouse (not publicly available) dataset called RP6K with 6500 retail products. This has over 1 Million images of retail products with their class tags. Just like real world, the number of images of these products in RP6K is long-tailed with some products having upto 4000 images, while many having lesser than 10 images. We finetune CLIP image encoder on this dataset using a novel mixture of techniques which yields a model that can be used for Zero shot classification on other datasets.
Grozi-120 is the first dataset used for evaluation [Merler et al., 2007]. It is a one shot classification dataset with one train image per product, which is packshot like, while test set images are from real world. It has 120 products.
@@
RP2K [Peng et al., 2020] is a dataset with 2388 retail products, with substantial number of both train and test images for each product. However, we still use the dataset in a few shot setting like Grozi120 and CAPG-GP. That is, while the entire test set will be used to test the performance of Zero Shot classification, only one image per class from train set will be used to calculate class representations for classification.
@@ 4 METHOD
Our work involves finetuning OpenCLIP image encoder on a retail product image dataset RP6K such that resultant image encoder can yield embeddings of product images useful for zero shot retail product image classification on other datasets. The image encoder after finetuning with RP6K is referred to as RetailKLIP. For zero shot classification, one image per product is taken and augmented and then passed to RetailKLIP to get embeddings for that product. This process is repeated for all products to create a set of products’ embeddings. To test an unseen image, its image embedding created using RetailKLIP is compared to product embeddings created earlier and the product with closest embedding is considered to be the output of RetailKLIP.
@@ 4.1 Finetuning OpenCLIP to RetailKLIP
We use the ViT-L OpenCLIP pretrained model to obtain RetailKLIP through finetuning. Specific Architecture used is ViT-L/14 trained on LAION-2B available on OpenCLIP’s Github [Ilharco et al., 2021]. We take the trick of using AdamW optimizer to finetune CLIP from [Kumar et al., 2022]. We also use differing learning rate with depth like in [Dong et al., 2022], but instead of layerwise rate decay, we use block wise rate decay. That is, within each block of ViT there is a common learning rate, but learning rate varies across blocks. RP6K is a very imbalanced dataset, hence we use ArcFace loss to finetune OpenCLIP. Specifically, ArcFace [Deng et al., 2019] implementation of PyTorch Metric Learning library [Musgrave et al., 2020] is used for the task. ArcFace is considered good for openset recognition and we use it because while the finetuning is being done on RP6K dataset, the model is used for Zero Shot classification on CAPG-GP, Grozi-120 and RP2K datasets, making it an openset task. ArcFace seems to also hold itself better against softmax on datasets with a long tail distribution according to previous work, infact just vanilla ArcFace finetuning is better than softmax finetuning with data balancing tricks [Zhang and Deng, 2020].
4.1.1 Techniques used for finetuning OpenCLIP on RP6K
CLIP models have proven to be hard to finetune. However, [Kumar et al., 2022] have proposed methods that can be used to finetune the CLIP vision-backbone for Image Recognition and other Computer Vision tasks. We use AdamW optimizer as is suggested. ArcFace [Deng et al., 2019] is used as loss function while finetuning.
4.1.2 Block wise Learning Rate Decay Trick
We also find like in the case of [Dong et al., 2022] that using a common learning rate for the entire model performs at much lower levels than using differing learning rates across the backbone. However, we replace the LLRD (Layerwise Learning Rate Decay) strategy of the work with Blockwise learning rate decay. LLRD starts with a learning rate for the top layer (classification head in case of cited work) and decreases learning rate for each layer by a constant multiplicative decay. In our work, we use a blockwise multiplicative decay strategy. The last ViT-L block (top block) starts with learning rate of 2e-4 and each previous block having a decreasing learning rate by a factor of 0.7. The blocks used here are blocks of transformers and other layers labelled in OpenCLIP’s code implementation of ViT.
4.1.3 Data Balancing
While ArcFace gives good results even without any data balancing while training, we use a custom data balancing approach to finetune on a steep imbalanced dataset like RP6K. While finetuning of RP6K, we create a held out validation set from RP6K dataset with equal number of test images from each product. Average accuracy on this validation set is used to determine optimal data balancing. [Zhang and Deng, 2020] takes 2 concepts into account while data balancing, depth, which is number of datapoints per class in train set and breadth, that is number of classes. We find that keeping breadth to maximum possible gets best average accuracy on the mentioned validation set. We test many values of depth to find the value which gets best avg accuracy on RP6K validation set. So maximum breadth and determined value of depth is used for finetuning. Classes are oversampled or undersampled accordingly with various augmentation tricks to get the balanced dataset.
@@ 4.2 Zero Shot classification using RetailKLIP
To use RetailKLIP as a Zero Shot classifier on a new dataset, one doesn’t need to finetune RetailKLIP. One image per product is taken from trainset and its augmentations are created. These augmentations of a product image are passed through RetailKLIP to get a set of embeddings for the product. Once we have embeddings for all products, we are ready to classify. An image of the test dataset is passed through RetailKLIP to get its embedding and compare it with existing products’ embeddings. The nearest embedding from train set is taken to be the result of classification. This is k-nearest-neighbors classification process with k=1.
@@ 5.1 RP2K results
RP2K dataset paper proposes full finetuning ResNet and other backbones on their dataset one category at a time to get upto 95% accuracy in some categories of their dataset. For some other harder categories, the accuracy can be lower than 90%. However, due to language barrier, our team has been unable to separate out categories and thus we have to use models on all categories combined, so this makes the problem harder for RetailKLIP. Also RetailKLIP takes just one image from train set to classify test images unlike full finetuning which uses the ample amount of images available. The accuracy for Zero Shot classification on RP2K is 87.7%. This is close to the accuracy full finetuning approach of [Peng et al., 2020] gets on the hardest categories.